OpenStudy (xapproachesinfinity):

any easy way to do this than half angle formulas \[\int_0^\pi\sin^4\theta \cos^2 \theta d\theta \]

OpenStudy (xapproachesinfinity):

the half angle formulas involve a lot of work, so i was wondering about another method

OpenStudy (solomonzelman):

I tried sin^4■ - sin^6■, but sin^6■ is too long to work out...

OpenStudy (solomonzelman):

let's see if I can find something good and quick

OpenStudy (xapproachesinfinity):

yeah too long steps, i got pretty tired lol

OpenStudy (xapproachesinfinity):

your aproach is the first thing i attempted then i moved to using half angles from the very first step but seeems really long

OpenStudy (xapproachesinfinity):

i think i will just try that first one again and hope i don't make a mistake

OpenStudy (solomonzelman):

it seems like one of the "rather long" problems....

OpenStudy (xapproachesinfinity):

i hate long ones i give up easily lol

OpenStudy (anonymous):

use those silly reduction formulas found in the back of the book

OpenStudy (solomonzelman):

everything that I do is becoming too long for a simple problem

OpenStudy (solomonzelman):

I haven't ever used them ...

OpenStudy (xapproachesinfinity):

neither do I, i think i will just go with that lol to reduce the trouble

OpenStudy (solomonzelman):

Yeah, I suppose.... (I mean I could solve it just by u-substitution in a year or two, but it isn't worth it, also I am doing another problem on my own time)

OpenStudy (xapproachesinfinity):

did over 60 problems my head is burning now lol there were many who were long

OpenStudy (solomonzelman):

(It is Σ 1/ln(Γ(kx)) for different k .... ) I will see if there are some open questions to do, but, I never liked integrals with a happy end but someone dies (integratable, but too long).

OpenStudy (solomonzelman):

using series for these integrals would be in a snap, but series isn't like really a complete answer though.

OpenStudy (solomonzelman):

In any case, good luck!! cu

Parth (parthkohli):

\[u = \tan x\]is a better substitution here, as I recall.

Parth (parthkohli):

ok no

OpenStudy (solomonzelman):

that looked scary...

OpenStudy (anonymous):

I don't know about easy, but if you're really intent on going around using the half-angle formulas, you can try various substitutions that exploit symmetry to rewrite the initial integral. \[\mathcal{I}=\int_0^\pi\sin^4t\cos^2t\,\mathrm{d}t\]Replacing \(t\) with \(\pi-t\), it's easy to see that \[\mathcal{I}=\int_0^\pi\cos^4t\sin^2t\,\mathrm{d}t\]Furthermore, replacing \(t\) with \(\dfrac{\pi}{2}-t\) gives \[\mathcal{I}=\int_{-\pi/2}^{\pi/2}\sin^4t\cos^2t\,\mathrm{d}t=\int_{-\pi/2}^{\pi/2}\cos^4t\sin^2t\,\mathrm{d}t\]Since both integrands are even, you get \[\frac{1}{2}\mathcal{I}=\int_0^{\pi/2}\sin^4t\cos^2t\,\mathrm{d}t=\int_0^{\pi/2}\cos^4t\sin^2t\,\mathrm{d}t\]which further tells you that \[\mathcal{I}=\int_0^{\pi/2}(\sin^4t+\cos^4t)\,\mathrm{d}t\]What can you do with this?

OpenStudy (anonymous):

DeMoivre's theorem could work in "simplifying" the current integrand. \[\begin{align*}\Re\left\{e^{4it}\right\}&=\Re\left\{(\cos t+i\sin t)^4\right\}\\[1ex] &=\cos^4t+\sin^4t-\cos^2t\sin^2t\\[1ex] &=\cos4t\end{align*}\]So you have \[\mathcal{I}=\int_0^{\pi/2}\left(\cos4t+\cos^2t\sin^2t\right)\,\mathrm{d}t\]

OpenStudy (anonymous):

From here, \[\cos^2t\sin^2t=\left(\frac{1}{2}\sin2t\right)^2=\frac{1}{4}\sin^22t\]Let \[\mathcal{J}=\int_0^{\pi/2}\sin^22t\,\mathrm{d}t\]so that \[\mathcal{I}=\int_0^{\pi/2}\cos4t\,\mathrm{d}t+\frac{1}{4}\mathcal{J}\]Replacing \(t\) with \(\dfrac{\pi}{4}-t\) and \(\dfrac{\pi}{2}-t\), you can show that \[\mathcal{J}=\int_0^{\pi/2}\sin^22t\,\mathrm{d}t=\int_0^{\pi/2}\cos^22t\,\mathrm{d}t\]so that \[\mathcal{J}=\frac{1}{2}\int_0^{\pi/2}\mathrm{d}t\]

OpenStudy (thomas5267):

That's even more work than the half-angle...

OpenStudy (xapproachesinfinity):

interesting approach @SithsAndGiggles